OpenStudy (anonymous):

In order to earn credit for the slope and intercepts, you must explicitly write each of these sentences, filling in the blank with the correct value for that problem: 1. The x-intercept is ___. 2. The y-intercept is ___. 3. The slope is ___. Merely listing numbers, or letters and numbers, is not sufficient to earn credit. 3. (5 points) y = -x + 4 4. (5 points) y = -3x - 8 5. (5 points) 4x = 9

OpenStudy (anonymous):

When x = 0, what axis must the point lie on? When y = 0, what axis must the point lie on? The slope can be found if the equation is in slope-intercept form. y = mx + b. m is the slope.

OpenStudy (anonymous):

so how do I figure out y=-x+4?

OpenStudy (anonymous):

Substitute 0 in for x and you will get the y-intercept. Substitute 0 in for y and solve for x and you will get the x-intercept. The value of m is the coefficient on the x variable. In this case, -1.

OpenStudy (anonymous):

so y+0=-0+4?

OpenStudy (anonymous):

y-intercept: y=-(0) + 4, solve for y x-intercept: (0)=-x + 4, solve for x and slope is the coefficient of x, in this case think of the "-" operator as multiplying x by (-1), thus your slope